# FocusMAE: Gallbladder Cancer Detection from Ultrasound Videos with Focused Masked Autoencoders

### v1

@@ Abstract
In recent years, automated Gallbladder Cancer (GBC) detection has gained the attention of researchers. Current state-of-the-art (SOTA) methodologies relying on ultrasound sonography (US) images exhibit limited generalization, emphasizing the need for transformative approaches. We observe that individual US frames may lack sufficient information to capture disease manifestation. This study advocates for a paradigm shift towards video-based GBC detection, leveraging the inherent advantages of spatiotemporal representations. Employing the Masked Autoencoder (MAE) for representation learning, we address shortcomings in conventional image-based methods. We propose a novel design called FocusMAE to systematically bias the selection of masking tokens from high-information regions, fostering a more refined representation of malignancy. Additionally, we contribute the most extensive US video dataset for GBC detection. We also note that, this is the first study on US video-based GBC detection. We validate the proposed methods on the curated dataset, and report a new state-of-the-art (SOTA) accuracy of 96.4% for the GBC detection problem, against an accuracy of 84% by current Image-based SOTA – GBCNet, and RadFormer, and 94.7% by Video-based SOTA – AdaMAE. We further demonstrate the generality of the proposed FocusMAE on a public CT-based Covid detection dataset, reporting an improvement in accuracy by 3.3% over current baselines. The source code and pretrained models are available at: https://github.com/sbasu276/FocusMAE.
@@ 2 Related Work
DNNs for GBC Detection. Despite numerous studies on DNNs for gallbladder-related diseases, only a few have explored AI-based detection of GBC [19]. Chang et. al[10] employed a UNet-based denoising to enhance the image quality of Low-Dose CT scans for characterizing GBC.In contrast, Basu et. al[5] introduced a CNN architecture called MS-SoP and a Gaussian blurring-based curriculum for efficient GBC detection in US images. Gupta et. al[20] further uses MS-SoP for a patient subgroup study on a large prospective patient cohort. [6] exploits a transformer-based dual-branch architecture with bag-of-word style feature embedding for accurate and explainable GBC detection. [21] investigates application of transformers for differentiation of GBC with xanthogranulomatous cholecystitis. [8], on the other hand, utilizes unsupervised contrastive learning to learn malignancy representations. Basu et. al[7] proposes DETR-based weakly supervised GBC detection. Gupta et. al[18] proposes a calibration metric and loss to calibrate the GBC detection models on small dataset. Despite the above studies, we observe a notable gap in the literature regarding models for video-based GBC detection from US videos. This gap in research motivates the current work.
@@ 3.3 Training
FocusMAE (Ours) Transformer 0.964$\pm$0.072 0.940$\pm$0.120 1.000$\pm$0.000
@@ 4.1 Curated US Video Dataset for GBC Detection
Annotation. The video labels in GBUSV are already provided. For our additional videos, we relied on the biopsy reports for labeling. Additionally, two radiologists with 2 and 10 years of expertise in abdominal ultrasound (US), were consulted to draw bounding boxes covering the entire GB and the adjacent liver parenchyma in the video frames.
@@
Dataset Splits. We report the 5-fold cross-validation metrics for the complete dataset for key experiments. The cross-validation splits were conducted on a patient-wise basis, ensuring that all videos of a particular patient appeared exclusively in either the training or the validation split during cross-validation.
@@ 6.2 Generality of the Proposed Method
Figure 5: Ablation study. We report the mean scores over 5-fold cross-validation for GBC detection. (a) Effect of varying the masking ratio ($\rho$) on accuracy. (b) Effect of varying the reconstruction loss - L1 vs. MSE - for SSL pretraining. Training with MSE yields 2.2% better accuracy. (c) Performance for different backbones. (d) Effect of varying the decoder depth.
@@ 6.3 Ablation Study
We performed the ablation study on the FocusMAE with the ViT backbone on the US Video dataset. We show the mean scores over the 5-fold cross-validation.
@@
Reconstruction Loss. We examined the effect of varying the reconstruction loss function in our study. We experimented with two variants: Mean Absolute Error or L1-loss, and Mean Squared Error (MSE). The results, shown in Fig. 4(b), indicate that using MSE loss during pretraining produces better performance in terms of accuracy. Models trained with MSE loss demonstrated a 2.2% higher mean accuracy compared to those trained with L1 loss.
@@ 7 Conclusion
This study addresses the limitations of current US image-based Gallbladder Cancer (GBC) detection techniques, emphasizing the need for a paradigm shift towards US video-based approaches. Our novel design, named FocusMAE, strategically biases masking token selection from high-information regions and learns quality representations of GB malignancy. FocusMAE achieves state-of-the-art results on US video-based GBC detection. We anticipate that our work will be valuable to the community, sparking interest in the challenging problem of GBC detection from US videos. Moreover, we showcase the generality of FocusMAE by applying it successfully to a public lung CT-based Covid detection task, demonstrating its applicability across two modalities and diseases. This suggests that FocusMAE could find broader use-cases in future, marking it a promising step towards versatile diagnostic solutions.
@@ Appendix C Visualization
Masking Ratio Decoder Depth Loss
MSE L1
Acc. Spec. Sens. Acc. Spec. Sens.
0.6 4 0.538$\pm$0.443 0.808$\pm$0.123 0.940$\pm$0.120 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
6 0.945$\pm$0.071 0.940$\pm$0.120 1.000$\pm$0.000 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
8 0.942$\pm$0.066 0.940$\pm$0.120 0.938$\pm$0.076 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
10 0.749$\pm$0.144 0.940$\pm$0.120 0.600$\pm$0.490 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
0.8 4 0.940$\pm$0.066 0.940$\pm$0.120 0.938$\pm$0.076 0.86$\pm$0.124 0.699$\pm$0.371 0.940$\pm$0.120
6 0.942$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076 0.943$\pm$0.065 1.000$\pm$0.000 0.938$\pm$0.076
8 0.942$\pm$0.066 0.940$\pm$0.120 0.938$\pm$0.076 0.978$\pm$0.027 1.000$\pm$0.000 0.938$\pm$0.076
10 0.952$\pm$0.070 0.940$\pm$0.120 0.967$\pm$0.067 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
0.95 4 0.810$\pm$0.124 0.940$\pm$0.120 0.538$\pm$0.443 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
6 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
8 0.811$\pm$0.122 0.940$\pm$0.120 0.538$\pm$0.443 0.943$\pm$0.065 0.940$\pm$0.120 0.938$\pm$0.076
10 0.964$\pm$0.072 0.940$\pm$0.120 1.000$\pm$0.000 0.943$\pm$0.065 0.094$\pm$0.120 0.938$\pm$0.076
Table S2: Hyperparameter tuning. We vary the Masking ratio, Decoder depth, and the loss function for a grid search. We report the 5-fold cross-validation (Mean±$\pm$±SD) accuracy, specificity, and sensitivity for FocusMAE in detecting GBC from the US.
Appendix D Hyperparameter Tuning
We have done rigorous grid searching for tuning the FocusMAE hyperparameters. We varied the masking ratio, decoder depth, and the loss function to select the optimal configuration over a grid search. Tab. S2 shows the results for different hyperparameter combinations.
@@ Appendix D Hyperparameter Tuning
VideoMAEv2 [49]
Vision transformer (ViT) backbone, with random masked auto-encoders. Masking in both encoder and decoders. All attention-based layers were trainable. ViT base model used for inference.
$3\times 16\times 224\times 224$
AdamW, LR = 7e-5, momentum = 0.999 ,weight decay = 0.1
4
30 epochs
TimeSformer [9]
Vision transformer based space time atention. Divided space-time attention configuration used. ViT base model used for inference.
$3\times 8\times 224\times 224$
SGD, LR = 0.005, weight decay=1e-4, momentum=0.9
8
25 epochs
VideoSwin [34]
Pretrained on ImageNet-1K. SwinTransformer3D based backbone. All layers were trainable
$3\times 8\times 224\times 224$
SGD, LR = 0.01, weight decay=1e-4, momentum=0.9
4
30 epochs
AdaMAE [4]
Vision transformer (ViT) backbone, with adaptively masked auto-encoders. The model embedding size provided by the authors is 768; we have pre-trained the 384 version to allow for a better fit to our data. Masking in only encoders. All attention-based layers were trainable. ViT base model used for inference
$3\times 16\times 224\times 224$
AdamW, LR = 1e-6, weight decay=0.9, momentum=0.99
2
10 epochs
VidTr [33]
Transformer-based video classification with separable attention. ViT-B backbone. All layers were trainable.
$3\times 16\times 224\times 224$
SGD, LR = 3e-4, weight decay=1e-5, momentum=0.9
2
40 epochs
@@
Appendix E Baseline Implementation Details

### v2

@@ Abstract
In recent years, automated Gallbladder Cancer (GBC) detection has gained the attention of researchers. Current state-of-the-art (SOTA) methodologies relying on ultrasound sonography (US) images exhibit limited generalization, emphasizing the need for transformative approaches. We observe that individual US frames may lack sufficient information to capture disease manifestation. This study advocates for a paradigm shift towards video-based GBC detection, leveraging the inherent advantages of spatiotemporal representations. Employing the Masked Autoencoder (MAE) for representation learning, we address shortcomings in conventional image-based methods. We propose a novel design called FocusMAE to systematically bias the selection of masking tokens from high-information regions, fostering a more refined representation of malignancy. Additionally, we contribute the most extensive US video dataset for GBC detection. We also note that, this is the first study on US video-based GBC detection. We validate the proposed methods on the curated dataset, and report a new SOTA accuracy of 96.4% for the GBC detection problem, against an accuracy of 84% by current Image-based SOTA – GBCNet and RadFormer, and 94.7% by Video-based SOTA – AdaMAE. We further demonstrate the generality of the proposed FocusMAE on a public CT-based Covid detection dataset, reporting an improvement in accuracy by 3.3% over current baselines. Project page with source code, trained models, and data is available at: https://gbc-iitd.github.io/focusmae.
@@ 2 Related Work
DNNs for GBC Detection. Despite numerous studies on DNNs for gallbladder-related diseases, only a few have explored AI-based detection of GBC [19]. Chang et. al[10] employed a UNet-based denoising to enhance the image quality of Low-Dose CT scans for characterizing GBC. In contrast, Basu et. al[5] introduced a CNN architecture called MS-SoP and a Gaussian blurring-based curriculum for efficient GBC detection in US images. Gupta et. al[20] further studied the performance of MS-SoP in classifying different sub-types of GBC on a large prospective patient cohort. Basu et. al[8] later utilized unsupervised contrastive learning to learn malignancy representations. On the other hand, [6] exploits a transformer-based dual-branch architecture for accurate and explainable GBC detection. [21] investigates application of transformers for differentiation of GBC with xanthogranulomatous cholecystitis. [7] further proposes DETR-based weakly supervised GBC detection. Gupta et. al[18] proposes a calibration metric and loss to calibrate the GBC detection models on small dataset. Despite the above studies, we observe a notable gap in the literature regarding models for video-based GBC detection from US videos. This gap in research motivates the current work.
@@ 3.3 Training
FocusMAE (Ours) Transformer 0.964$\pm$0.047 0.910$\pm$0.117 1.000$\pm$0.000
@@ 4.1 Curated US Video Dataset for GBC Detection
Annotation. The video labels in GBUSV are already provided. For our additional videos, we relied on the biopsy reports for labeling. Additionally, two radiologists with 2 and 10 years of expertise in abdominal ultrasound (US), were consulted to draw bounding boxes covering the entire GB and the adjacent liver parenchyma in one frame in each video.
@@
Dataset Splits. We report the 5-fold cross-validation metrics over the complete dataset for key experiments. The cross-validation splits were conducted on a patient-wise basis, ensuring that all videos of a particular patient appeared exclusively in either the training or the validation split during cross-validation.
@@ 6.2 Generality of the Proposed Method
Figure 5: Ablation study. We report the mean scores over 5-fold cross-validation for GBC detection. (a) Effect of varying the masking ratio ($\rho$) on accuracy. (b) Effect of varying the reconstruction loss - L1 vs. MSE - for SSL pretraining. Training with MSE yields 2.1% better accuracy. (c) Performance for different backbones. (d) Effect of varying the decoder depth.
@@ 6.3 Ablation Study
We performed the ablation study on the FocusMAE with the ViT backbone on the US Video dataset.
@@
Reconstruction Loss. We examined the effect of varying the reconstruction loss function in our study. We experimented with two variants: Mean Absolute Error or L1-loss, and Mean Squared Error (MSE). The results, shown in Fig. 5(b), indicate that using MSE loss during pretraining produces slightly better performance in terms of accuracy. Models trained with MSE loss demonstrated 2.1% higher mean accuracy compared to those trained with L1 loss.
@@ 7 Conclusion
This study addresses the limitations of current US image-based GBC detection techniques, emphasizing the need for a paradigm shift towards US video-based approaches. Our novel design, named FocusMAE, strategically biases masking token selection from high-information regions and learns quality representations of GB malignancy. FocusMAE achieves state-of-the-art results on US video-based GBC detection. We hope that our work will spark interest in the challenging problem of GBC detection from US videos. Moreover, we showcase the generality of FocusMAE by applying it successfully to a public lung CT-based Covid detection task, demonstrating its applicability across two modalities and diseases. This suggests that FocusMAE could find broader use-cases in future, marking it a promising step towards versatile diagnostic solutions.
Acknowledgments. Authors thank Dr. Shravya Singh and Dr. Ruby Siddiqui for data annotation, and Dr. Pratyaksha Rana for reading the cases. This work was partially supported by the CSE Research Acceleration Fund of IIT Delhi.
@@ Appendix C Visualization
VideoMAEv2 [49] Vision transformer (ViT) backbone, with random masked auto-encoders. Masking in both encoder and decoders. All attention-based layers were trainable. ViT base model used for inference. $3\times 16\times 224\times 224$ AdamW, LR = 7e-5, momentum = 0.999 ,weight decay = 0.1 4 30 epochs
TimeSformer [9] Vision transformer based space time atention. Divided space-time attention configuration used. ViT base model used for inference. $3\times 8\times 224\times 224$ SGD, LR = 0.005, weight decay=1e-4, momentum=0.9 8 25 epochs
VideoSwin [34] Pretrained on ImageNet-1K. SwinTransformer3D based backbone. All layers were trainable $3\times 8\times 224\times 224$ SGD, LR = 0.01, weight decay=1e-4, momentum=0.9 4 30 epochs
AdaMAE [4] Vision transformer (ViT) backbone, with adaptively masked auto-encoders. The model embedding size provided by the authors is 768; we have pre-trained the 384 version to allow for a better fit to our data. Masking in only encoders. All attention-based layers were trainable. ViT base model used for inference $3\times 16\times 224\times 224$ AdamW, LR = 1e-6, weight decay=0.9, momentum=0.99 2 10 epochs
VidTr [33] Transformer-based video classification with separable attention. ViT-B backbone. All layers were trainable. $3\times 16\times 224\times 224$ SGD, LR = 3e-4, weight decay=1e-5, momentum=0.9 2 40 epochs
@@
Appendix D Baseline Implementation Details
@@
Appendix E Clip-level Statistics
We have a total of 484 clips sub-sampled from the 91 videos at the fine-tuning stage. Out of these, 320 clips were from the malignant videos, and contain the malignant label as per the positive biopsy reports. All clips of a malignant video is given the malignant label. Radiologists identified 199 clips out of these 320 to be malignant. At a frame-level, radiologists identified 3212 frames exhibiting signs of malignancy.